# A discrete model for Gell-Mann matrices

### 22 Jan 2024 v1

@@ 6.1. Summary
The suggestion that physics might be fundamentally discrete, and that the apparently continuous behaviour we often observe is an illusion, goes back a long way. In the context of quantum mechanics, this suggestion was made by Einstein [24] as early as 1935, but no discrete theory was found to rival the continuous models based on the Schrödinger equation and then the Dirac equation. The Standard Model of Particle Physics, largely completed by the mid 1970s, is entirely based on continuous groups, so much so that the emergence of discrete experimental observations is still not explained in a totally convincing way [25, 26, 27].
Theories beyond the Standard Model are almost without exception continuous theories. A few lone voices, such as ’t Hooft [28], still speak out in favour of Einstein’s dream of a discrete foundation, even if this has to be at the Planck scale. But a convincing discrete model is still lacking, and few people are seriously looking for one. The ones I have put forward in [9, 29, 30, 31, 10, 32] all suffer from drawbacks of various kinds, most notably the lack of any convincing analogue of the colour $SU(3)$, which is an essential part of the Standard Model.
Nevertheless, it is a mathematical fact that finite groups are much more complicated than Lie groups, and quite small finite groups can model properties that cannot be modelled with Lie groups at all. In this paper, therefore, I have tried to remedy this defect by explicitly proposing a finite analogue or precursor of $SU(3)$. At a combinatorial (qualitative) level it has many desirable properties that mirror known experimental properties of elementary particles, in particular the chirality of the weak interaction, and the mixing of electromagnetism and the weak and strong nuclear forces. Quantitative analysis of masses, mixing angles and probability amplitudes is beyond the scope of this paper.
@@ 6.2. Further work
The proposed unification of Gell-Mann and Pauli matrices using a 3-dimensional complex reflection group really does nothing more than provide a discrete foundation for the Standard Model of Particle Physics as it currently stands. Nevertheless, it permits new methods of calculation, using finite geometries, which have been shown to be effective in calculating four of the mixing angles from fundamental mass parameters. A priority for further work is therefore to try to find similar calculations for the remaining five mixing angles.
Further unification with the Dirac matrices cannot be done in the Standard Model, since it is impossible to arrange for the Lorentz group to commute with the whole of the gauge group in this case. What this means is that a putative unification with Dirac matrices provides a possible mechanism for particle physics to influence the shape of spacetime, and hence couple to gravity. Investigating this possibility further is a priority for future work.
@@
In order to gauge $15$ masses we need an $8$-dimensional complex representation, or a $4$-dimensional quaternionic representation. Since $SL(4,\mathbb{H})$ is a real form of $SU(8)$, of Lie type $A_{7}$, and since $A_{7}$ is contained in both $E_{7}$ and $D_{8}$, there are a number of ways to embed such a gauge group in (complex) $E_{8}$. It may even be possible to embed $SL(4,\mathbb{H})$ in a real form of $E_{8}$. In the $E_{7}$ case, however, there are no spinor representations of $SL(4,\mathbb{H})$ in $E_{8}$, which seems to rule this one out. In the $D_{8}$ case, the spinors split into $4+28$ quaternionic dimensions, which may be viable as a splitting into neutrinos versus the charged fundamental fermions. Even so, the model would still be quite unlike the various $E_{8}$ models [13, 33, 34, 35] that have been proposed over a number of years.

### 25 Jan 2024 v2

@@ 10.1. Summary
The suggestion that physics might be fundamentally discrete, and that the apparently continuous behaviour we often observe is an illusion, goes back a long way. In the context of quantum mechanics, this suggestion was made by Einstein [47] as early as 1935, but no discrete theory was found to rival the continuous models based on the Schrödinger equation and then the Dirac equation. The Standard Model of Particle Physics, largely completed by the mid 1970s, is entirely based on continuous groups, so that the emergence of discrete experimental observations is still not explained in a totally convincing way [48, 49, 50].
Theories beyond the Standard Model are almost without exception continuous theories. A few lone voices, such as ’t Hooft [51], still speak out in favour of Einstein’s dream of a discrete foundation, even if this has to be at the Planck scale. But a convincing discrete model is still lacking, and few people are seriously looking for one. The ones I have put forward in [9, 52, 53, 54, 10, 55] all suffer from drawbacks of various kinds, most notably the lack of any convincing analogue of the colour $SU(3)$, which is an essential and very successful part of the Standard Model. Nevertheless, it is a mathematical fact that finite groups are much more complicated than Lie groups, and quite small finite groups can model properties that cannot be modelled with Lie groups at all.
In this paper, therefore, I have tried to remedy this defect by explicitly proposing a finite analogue or precursor of $SU(3)$. At a combinatorial (qualitative) level it has many desirable properties that mirror known experimental properties of elementary particles, in particular the chirality of the weak interaction, the mixing of electromagnetism and the weak and strong nuclear forces, and neutrino oscillations. Quantitative analysis of masses, mixing angles and probability amplitudes is, however, beyond the scope of this paper, as it requires detailed investigations of coordinates for the representations, that may even (counter-intuitively) depend on the large-scale gravitational environment.
@@ 10.2. Further work
The proposed unification of Gell-Mann and Pauli matrices using a 3-dimensional complex reflection group really does little more than provide a discrete foundation for the Standard Model of Particle Physics as it currently stands. Nevertheless, it permits new methods of calculation, using finite geometries, which have been shown to be effective in calculating four of the mixing angles from fundamental mass parameters. A priority for further work is therefore to try to find similar calculations for the remaining five mixing angles.
Further unification with the Dirac matrices cannot be done in the Standard Model, since it is impossible to arrange for the Lorentz group to commute with the whole of the gauge group in this case. Nevertheless, the non-commutation is very subtle, and only really affects properties of neutrinos and gluons, as well as the Higgs field, so that any differences from a commuting theory are hard to detect. What this means is that a unification with Dirac matrices provides a possible mechanism for particle physics to influence the shape of spacetime, and hence couple to gravity. In other words, the proposed unification using a quaternionic reflection group provides a potential foundation for a quantum theory of gravity.
Investigating this possibility further is a priority for future work. The appropriate compact gauge group here is $Sp(4)$, of dimension $36$, but it is also possible to gauge the theory with the noncompact group $SL(4,\mathbb{H})$ of dimension $63$. In this case, there is a subgroup $SL(4,\mathbb{R})$ that might be used for general covariance, in order to reduce any such putative quantum theory of gravity to General Relativity in the appropriate limit. Since $SL(4,\mathbb{R})$ contains no quaternions, it cannot model the three generations of electrons, which suggests that it is the (quantum and/or classical) gravitational properties of electrons [56] that need to be investigated in case of any apparent deviations [57, 58] of gravity from General Relativity.
@@
Since $SL(4,\mathbb{H})$ is a real form of $SU(8)$, of Lie type $A_{7}$, and since $A_{7}$ is contained in both $E_{7}$ and $D_{8}$, there are at least two ways to embed this gauge group in (complex) $E_{8}$. It may even be possible to embed it in a real form of $E_{8}$. In the $E_{7}$ case, there are no spinor representations of $SL(4,\mathbb{H})$ in $E_{8}$, which seems to rule this one out [59, 60]. In the $D_{8}$ case, the spinors split into $4+28$ quaternionic dimensions, which may be viable as a splitting into neutrinos versus the charged fundamental fermions. Even so, the model would still be quite unlike the various $E_{8}$ models [16, 45, 61, 62] that have been proposed over a number of years.